# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 19.4.3 Image Style Transfer
Image style transfer is another fascinating application of GANs in which the goal is to transform an image from one style to another. For instance, using GANs, a photograph can be converted into a painting, or an image can be transformed from one artistic style to another. One of the most popular GAN architectures for style transfer is CycleGAN, which enables unpaired image-to-image translation, meaning it does not require paired training data [54].
@@
CycleGAN works by learning two sets of mappings: one that translates images from domain A to domain B (e.g., from photos to paintings) and another that translates images from domain B back to domain A (e.g., from paintings to photos). The key innovation in CycleGAN is the use of cycle consistency loss, which ensures that after transforming an image from one domain to another and back again, the image remains unchanged [222].
@@
2 class CycleGANGenerator(nn.Module):
@@
5 self.main = nn.Sequential(
6 nn.Conv2d(3, 64, kernel_size=7, padding=3),
7 nn.ReLU(inplace=True),
8 nn.Conv2d(64, 128, kernel_size=3, stride=2, padding=1),
9 nn.ReLU(inplace=True),
10 nn.Conv2d(128, 256, kernel_size=3, stride=2, padding=1),
11 nn.ReLU(inplace=True),
12 nn.ConvTranspose2d(256, 128, kernel_size=3, stride=2, padding=1, output_padding=1),
13 nn.ReLU(inplace=True),
14 nn.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1),
15 nn.ReLU(inplace=True),
16 nn.Conv2d(64, 3, kernel_size=7, padding=3)
@@
  • Artistic Style Transfer: Convert photographs into artistic paintings or transform images between different artistic styles [212].
  • Domain Adaptation: Translate images from one domain to another (e.g., converting night-time images to day-time images) [190].
@@
Autoencoders are a type of neural network designed for unsupervised learning tasks, where the goal is to compress and then reconstruct data as accurately as possible. In image processing, autoencoders can be used for tasks like image denoising, anomaly detection, and even generative image modeling. The core idea of autoencoders is to learn a lower-dimensional representation (or encoding) of the input image and use it to reconstruct the original image through a decoding process [71]. This chapter will explore the fundamental structure of autoencoders and introduce Variational Autoencoders (VAEs), which add a probabilistic component to the encoding process.
@@ 20.1.1 Encoder and Decoder
[…]
@@
5class Autoencoder(nn.Module):
@@
10 self.encoder = nn.Sequential(
11 nn.Conv2d(1, 16, kernel_size=3, stride=2, padding=1), # Output: 16x14x14
12 nn.ReLU(),
13 nn.Conv2d(16, 32, kernel_size=3, stride=2, padding=1), # Output: 32x7x7
14 nn.ReLU()
@@
18 self.decoder = nn.Sequential(
19 nn.ConvTranspose2d(32, 16, kernel_size=3, stride=2, padding=1, output_padding=1), # Output: 16x14x14
20 nn.ReLU(),
21 nn.ConvTranspose2d(16, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # Output: 1x28x28
22 nn.Sigmoid() # Use Sigmoid to get pixel values between [0, 1]
@@ 20.2 Variational Autoencoders (VAE)
Variational Autoencoders (VAEs) are a type of autoencoder that introduce a probabilistic approach to the latent space representation. Unlike standard autoencoders that learn a fixed deterministic encoding for each input, VAEs learn a distribution over the latent space, allowing for more diverse and smooth image generation [87]. VAEs are particularly useful for generating new images that are similar to those in the training set, making them popular in tasks like image synthesis.
@@
6 self.encoder = nn.Sequential(
7 nn.Conv2d(1, 16, kernel_size=3, stride=2, padding=1),
8 nn.ReLU(),
9 nn.Conv2d(16, 32, kernel_size=3, stride=2, padding=1),
10 nn.ReLU()
@@
12 self.fc_mu = nn.Linear(32*7*7, 20) # Mean vector
13 self.fc_logvar = nn.Linear(32*7*7, 20) # Log-variance vector
@@
16 self.decoder_fc = nn.Linear(20, 32*7*7)
17 self.decoder = nn.Sequential(
18 nn.ConvTranspose2d(32, 16, kernel_size=3, stride=2, padding=1, output_padding=1),
19 nn.ReLU(),
20 nn.ConvTranspose2d(16, 1, kernel_size=3, stride=2, padding=1, output_padding=1),
21 nn.Sigmoid() # Output pixel values in the range [0, 1]
@@
47 recon_loss = nn.functional.binary_cross_entropy(recon_x, x, reduction=’sum’)
@@ 20.4 Applications of Autoencoders in Image Processing
Autoencoders are a type of neural network designed to learn efficient representations of input data, often used for unsupervised learning tasks. They consist of two main components: an encoder that compresses the input into a lower-dimensional latent space and a decoder that reconstructs the input from this compressed representation. Autoencoders are particularly useful in image processing tasks such as image denoising, compression, and reconstruction. In this section, we will explore how autoencoders are applied in these areas, providing detailed explanations and examples [106].
@@ 20.4.1 Image Denoising
Denoising Autoencoders (DAEs) are a variant of autoencoders specifically designed to remove noise from images. The key idea is to train the autoencoder to map noisy inputs to clean outputs. The encoder compresses the noisy input, and the decoder learns to reconstruct the clean version of the image by learning noise-invariant features [202].
@@
8 class DenoisingAutoencoder(nn.Module):
@@
12 self.encoder = nn.Sequential(
13 nn.Conv2d(1, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 14, 14)
14 nn.ReLU(True),
15 nn.Conv2d(64, 128, kernel_size=3, stride=2, padding=1), # (batch_size, 128, 7, 7)
16 nn.ReLU(True)
@@
19 self.decoder = nn.Sequential(
20 nn.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 64, 14, 14)
21 nn.ReLU(True),
22 nn.ConvTranspose2d(64, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
23 nn.Sigmoid() # Output between 0 and 1
@@ 20.4.2 Image Compression
1 class CompressionAutoencoder(nn.Module):
@@
5 self.encoder = nn.Sequential(
6 nn.Conv2d(1, 32, kernel_size=3, stride=2, padding=1), # (batch_size, 32, 14, 14)
7 nn.ReLU(True),
8 nn.Conv2d(32, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 7, 7)
9 nn.ReLU(True),
10 nn.Conv2d(64, 128, kernel_size=7) # (batch_size, 128, 1, 1) compressed representation
@@
13 self.decoder = nn.Sequential(
14 nn.ConvTranspose2d(128, 64, kernel_size=7), # (batch_size, 64, 7, 7)
15 nn.ReLU(True),
16 nn.ConvTranspose2d(64, 32, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 32, 14, 14)
17 nn.ReLU(True),
18 nn.ConvTranspose2d(32, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
19 nn.Sigmoid()
@@ 20.4.3 Image Reconstruction and Inpainting
Image inpainting refers to the task of reconstructing missing or damaged parts of an image. Autoencoders can be used for inpainting by learning how to generate plausible content for the missing regions based on the surrounding pixels. This is achieved by training the autoencoder on images where certain parts are masked or corrupted, and the autoencoder learns to reconstruct the original, complete image [197].
@@
1 class InpaintingAutoencoder(nn.Module):
@@
5 self.encoder = nn.Sequential(
6 nn.Conv2d(1, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 14, 14)
7 nn.ReLU(True),
8 nn.Conv2d(64, 128, kernel_size=3, stride=2, padding=1), # (batch_size, 128, 7, 7)
9 nn.ReLU(True)
@@
12 self.decoder = nn.Sequential(
13 nn.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 64, 14, 14)
14 nn.ReLU(True),
15 nn.ConvTranspose2d(64, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
16 nn.Sigmoid()
@@
Image super-resolution and deblurring are two important tasks in image processing that focus on improving image quality. Super-resolution involves upscaling low-resolution images to generate high-resolution counterparts, while deblurring removes blur from images to restore sharpness. Traditional methods for these tasks often struggle to generate fine details or high-quality results [5]. However, deep learning techniques have led to significant advancements in both fields. In this chapter, we explore deep learning methods such as SRGAN, ESRGAN, VDSR, and EDSR for image super-resolution.
@@ 21.1 Image Super-Resolution
Image super-resolution aims to enhance the resolution of images by generating high-frequency details that were lost in the downscaling process. Deep learning models have proven highly effective in this task by learning complex mappings from low-resolution (LR) to high-resolution (HR) images [45].
@@ 21.1.1 SRGAN (Super-Resolution GAN)
Super-Resolution GAN (SRGAN) is a deep learning model that uses the adversarial training framework of GANs to produce realistic, high-resolution images from low-resolution inputs. The core of SRGAN consists of two networks: the generator, which generates high-resolution images, and the discriminator, which distinguishes between real and generated high-resolution images [207].
@@
The generator in SRGAN is a convolutional neural network (CNN) that takes a low-resolution image as input and outputs a high-resolution image. It is designed to predict high-frequency details that are missing in the low-resolution image. The generator is trained to ”fool” the discriminator by producing images that appear realistic [99].
@@
4class ResidualBlock(nn.Module):
@@
7 self.block = nn.Sequential(
8 nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1),
9 nn.BatchNorm2d(num_channels),
10 nn.PReLU(),
11 nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1),
12 nn.BatchNorm2d(num_channels)
@@
18class SRGAN_Generator(nn.Module):
@@
23 self.conv1 = nn.Conv2d(3, 64, kernel_size=9, padding=4)
24 self.prelu = nn.PReLU()
@@
27 self.residual_blocks = nn.Sequential(
@@
32 self.upsample = nn.Sequential(
33 nn.Conv2d(64, 256, kernel_size=3, padding=1),
34 nn.PixelShuffle(2), # Upscales by a factor of 2
35 nn.PReLU(),
36 nn.Conv2d(64, 256, kernel_size=3, padding=1),
37 nn.PixelShuffle(2), # Upscales by another factor of 2
38 nn.PReLU()
@@
42 self.conv2 = nn.Conv2d(64, 3, kernel_size=9, padding=4)
43 self.tanh = nn.Tanh() # Output in range [-1, 1]
@@
  • The output of the generator is a high-resolution image with pixel values in the range $[-1,1]$, produced by the Tanh activation function.

### v2

@@ 19.4.3 Image Style Transfer
Image style transfer is another fascinating application of GANs in which the goal is to transform an image from one style to another. For instance, using GANs, a photograph can be converted into a painting, or an image can be transformed from one artistic style to another. One of the most popular GAN architectures for style transfer is CycleGAN, which enables unpaired image-to-image translation, meaning it does not require paired training data [gatys2016image].
@@
CycleGAN works by learning two sets of mappings: one that translates images from domain A to domain B (e.g., from photos to paintings) and another that translates images from domain B back to domain A (e.g., from paintings to photos). The key innovation in CycleGAN is the use of cycle consistency loss, which ensures that after transforming an image from one domain to another and back again, the image remains unchanged [zhu2017unpaired].
@@
2 class CycleGANGenerator($n$.Module):
@@
5 self.main = $n$.Sequential(
6 $n$.Conv2d(3, 64, kernel_size=7, padding=3),
7 $n$.ReLU(inplace=True),
8 $n$.Conv2d(64, 128, kernel_size=3, stride=2, padding=1),
9 $n$.ReLU(inplace=True),
10 $n$.Conv2d(128, 256, kernel_size=3, stride=2, padding=1),
11 $n$.ReLU(inplace=True),
12 $n$.ConvTranspose2d(256, 128, kernel_size=3, stride=2, padding=1, output_padding=1),
13 $n$.ReLU(inplace=True),
14 $n$.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1),
15 $n$.ReLU(inplace=True),
16 $n$.Conv2d(64, 3, kernel_size=7, padding=3)
@@
  • Artistic Style Transfer: Convert photographs into artistic paintings or transform images between different artistic styles [xu2021drb].
  • Domain Adaptation: Translate images from one domain to another (e.g., converting night-time images to day-time images) [sun2019see].
@@
Autoencoders are a type of neural network designed for unsupervised learning tasks, where the goal is to compress and then reconstruct data as accurately as possible. In image processing, autoencoders can be used for tasks like image denoising, anomaly detection, and even generative image modeling. The core idea of autoencoders is to learn a lower-dimensional representation (or encoding) of the input image and use it to reconstruct the original image through a decoding process [hinton2006reducing]. This chapter will explore the fundamental structure of autoencoders and introduce Variational Autoencoders (VAEs), which add a probabilistic component to the encoding process.
@@ 20.1.1 Encoder and Decoder
$z=f_{\text{enc}}(x)\quad\text{(encoder)}$
$\hat{x}=f_{\text{dec}}(z)\quad\text{(decoder)}$
@@
5class Autoencoder($n$.Module):
@@
10 self.encoder = $n$.Sequential(
11 $n$.Conv2d(1, 16, kernel_size=3, stride=2, padding=1), # Output: 16x14x14
12 $n$.ReLU(),
13 $n$.Conv2d(16, 32, kernel_size=3, stride=2, padding=1), # Output: 32x7x7
14 $n$.ReLU()
@@
18 self.decoder = $n$.Sequential(
19 $n$.ConvTranspose2d(32, 16, kernel_size=3, stride=2, padding=1, output_padding=1), # Output: 16x14x14
20 $n$.ReLU(),
21 $n$.ConvTranspose2d(16, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # Output: 1x28x28
22 $n$.Sigmoid() # Use Sigmoid to get pixel values between [0, 1]
@@ 20.2 Variational Autoencoders (VAE)
Variational Autoencoders (VAEs) are a type of autoencoder that introduce a probabilistic approach to the latent space representation. Unlike standard autoencoders that learn a fixed deterministic encoding for each input, VAEs learn a distribution over the latent space, allowing for more diverse and smooth image generation [kingma2013auto]. VAEs are particularly useful for generating new images that are similar to those in the training set, making them popular in tasks like image synthesis.
@@
6 self.encoder = $n$.Sequential(
7 $n$.Conv2d(1, 16, kernel_size=3, stride=2, padding=1),
8 $n$.ReLU(),
9 $n$.Conv2d(16, 32, kernel_size=3, stride=2, padding=1),
10 $n$.ReLU()
@@
12 self.fc_mu = $n$.Linear(32*7*7, 20) # Mean vector
13 self.fc_logvar = $n$.Linear(32*7*7, 20) # Log-variance vector
@@
16 self.decoder_fc = $n$.Linear(20, 32*7*7)
17 self.decoder = $n$.Sequential(
18 $n$.ConvTranspose2d(32, 16, kernel_size=3, stride=2, padding=1, output_padding=1),
19 $n$.ReLU(),
20 $n$.ConvTranspose2d(16, 1, kernel_size=3, stride=2, padding=1, output_padding=1),
21 $n$.Sigmoid() # Output pixel values in the range [0, 1]
@@
47 recon_loss = $n$.functional.binary_cross_entropy(recon_x, x, reduction=’sum’)
@@ 20.4 Applications of Autoencoders in Image Processing
Autoencoders are a type of neural network designed to learn efficient representations of input data, often used for unsupervised learning tasks. They consist of two main components: an encoder that compresses the input into a lower-dimensional latent space and a decoder that reconstructs the input from this compressed representation. Autoencoders are particularly useful in image processing tasks such as image denoising, compression, and reconstruction. In this section, we will explore how autoencoders are applied in these areas, providing detailed explanations and examples [li2023comprehensive].
@@ 20.4.1 Image Denoising
Denoising Autoencoders (DAEs) are a variant of autoencoders specifically designed to remove noise from images. The key idea is to train the autoencoder to map noisy inputs to clean outputs. The encoder compresses the noisy input, and the decoder learns to reconstruct the clean version of the image by learning noise-invariant features [vincent2008extracting].
@@
8 class DenoisingAutoencoder($n$.Module):
@@
12 self.encoder = $n$.Sequential(
13 $n$.Conv2d(1, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 14, 14)
14 $n$.ReLU(True),
15 $n$.Conv2d(64, 128, kernel_size=3, stride=2, padding=1), # (batch_size, 128, 7, 7)
16 $n$.ReLU(True)
@@
19 self.decoder = $n$.Sequential(
20 $n$.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 64, 14, 14)
21 $n$.ReLU(True),
22 $n$.ConvTranspose2d(64, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
23 $n$.Sigmoid() # Output between 0 and 1
@@ 20.4.2 Image Compression
1 class CompressionAutoencoder($n$.Module):
@@
5 self.encoder = $n$.Sequential(
6 $n$.Conv2d(1, 32, kernel_size=3, stride=2, padding=1), # (batch_size, 32, 14, 14)
7 $n$.ReLU(True),
8 $n$.Conv2d(32, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 7, 7)
9 $n$.ReLU(True),
10 $n$.Conv2d(64, 128, kernel_size=7) # (batch_size, 128, 1, 1) compressed representation
@@
13 self.decoder = $n$.Sequential(
14 $n$.ConvTranspose2d(128, 64, kernel_size=7), # (batch_size, 64, 7, 7)
15 $n$.ReLU(True),
16 $n$.ConvTranspose2d(64, 32, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 32, 14, 14)
17 $n$.ReLU(True),
18 $n$.ConvTranspose2d(32, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
19 $n$.Sigmoid()
@@ 20.4.3 Image Reconstruction and Inpainting
Image inpainting refers to the task of reconstructing missing or damaged parts of an image. Autoencoders can be used for inpainting by learning how to generate plausible content for the missing regions based on the surrounding pixels. This is achieved by training the autoencoder on images where certain parts are masked or corrupted, and the autoencoder learns to reconstruct the original, complete image [tu2019facial].
@@
1 class InpaintingAutoencoder($n$.Module):
@@
5 self.encoder = $n$.Sequential(
6 $n$.Conv2d(1, 64, kernel_size=3, stride=2, padding=1), # (batch_size, 64, 14, 14)
7 $n$.ReLU(True),
8 $n$.Conv2d(64, 128, kernel_size=3, stride=2, padding=1), # (batch_size, 128, 7, 7)
9 $n$.ReLU(True)
@@
12 self.decoder = $n$.Sequential(
13 $n$.ConvTranspose2d(128, 64, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 64, 14, 14)
14 $n$.ReLU(True),
15 $n$.ConvTranspose2d(64, 1, kernel_size=3, stride=2, padding=1, output_padding=1), # (batch_size, 1, 28, 28)
16 $n$.Sigmoid()
@@
Image super-resolution and deblurring are two important tasks in image processing that focus on improving image quality. Super-resolution involves upscaling low-resolution images to generate high-resolution counterparts, while deblurring removes blur from images to restore sharpness. Traditional methods for these tasks often struggle to generate fine details or high-quality results [albluwi2018image]. However, deep learning techniques have led to significant advancements in both fields. In this chapter, we explore deep learning methods such as SRGAN, ESRGAN, VDSR, and EDSR for image super-resolution.
@@ 21.1 Image Super-Resolution
Image super-resolution aims to enhance the resolution of images by generating high-frequency details that were lost in the downscaling process. Deep learning models have proven highly effective in this task by learning complex mappings from low-resolution (LR) to high-resolution (HR) images [dong2015image].
@@ 21.1.1 SRGAN (Super-Resolution GAN)
Super-Resolution GAN (SRGAN) is a deep learning model that uses the adversarial training framework of GANs to produce realistic, high-resolution images from low-resolution inputs. The core of SRGAN consists of two networks: the generator, which generates high-resolution images, and the discriminator, which distinguishes between real and generated high-resolution images [wang2023review].
@@
The generator in SRGAN is a convolutional neural network (CNN) that takes a low-resolution image as input and outputs a high-resolution image. It is designed to predict high-frequency details that are missing in the low-resolution image. The generator is trained to ”fool” the discriminator by producing images that appear realistic [ledig2017photo].
@@
4class ResidualBlock($n$.Module):
@@
7 self.block = $n$.Sequential(
8 $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1),
9 $n$.BatchNorm2d(num_channels),
10 $n$.PReLU(),
11 $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1),
12 $n$.BatchNorm2d(num_channels)
@@
18class SRGAN_Generator($n$.Module):
@@
23 self.conv1 = $n$.Conv2d(3, 64, kernel_size=9, padding=4)
24 self.prelu = $n$.PReLU()
@@
27 self.residual_blocks = $n$.Sequential(
@@
32 self.upsample = $n$.Sequential(
33 $n$.Conv2d(64, 256, kernel_size=3, padding=1),
34 $n$.PixelShuffle(2), # Upscales by a factor of 2
35 $n$.PReLU(),
36 $n$.Conv2d(64, 256, kernel_size=3, padding=1),
37 $n$.PixelShuffle(2), # Upscales by another factor of 2
38 $n$.PReLU()
@@
42 self.conv2 = $n$.Conv2d(64, 3, kernel_size=9, padding=4)
43 self.tanh = $n$.Tanh() # Output in range [-1, 1]
@@
  • The output of the generator is a high-resolution image with pixel values in the range [1,1][-1,1], produced by the Tanh activation function.